# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 2.3.2 Function similarity for general convex functions
Assume that for all $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$. We say that the set of functions $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar if there exists a subgradient oracle $\nabla F:\overline{\mathcal{X}}\to\mathbb{R}^{md}$, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v}))$, such that for each $s\in[S]$,
@@
$a_{1}^{2}\geq\sup_{x\in\mathcal{X}}\sum_{v\in V}\|f^{\prime}_{v}(x)-\frac{1}{m}\sum_{v^{\prime}\in V}f_{v^{\prime}}^{\prime}(x)\|^{2}.$
@@
Comparisons with existing notions of function similarity. [pmlr-v119-karimireddy20a] proposes the bounded gradient dissimilarity for differentiable convex objectives, which coincides with our Definition 2.1 when $S=1$ and when the objectives are differentiable. For twice differentiable objectives, function similarity is also defined in terms of differences in Hessians, i.e. $\|\nabla^{2}f_{v}-\nabla^{2}f_{v^{\prime}}\|$ [pmlr-v151-tian22b, konecny2016federatedoptimizationdistributedmachine, ArjevaniShamir2015Communication, pmlr-v119-karimireddy20a]. For general convex functions which could be non-differentiable, [ArjevaniShamir2015Communication] informally defines it ($\delta$-relatedness in their terminology) as the condition that “subgradients of local functions are at most $\delta$-different from each other”. Our Definition 2.1 formalizes this idea, and extend it to the case where $S>1$.
@@ 2.3.3 Requirements on $R_{s}$ without orthogonality
The above Lemma 2.3 imposes orthogonality assumptions on $\{K_{s}\}_{s\in[S]}$. In the more general case where such assumptions do not hold, one can always view ($\mathcal{P}_{s}$) as a problem with only $1$ block, with $K$ and $R$ as the corresponding operator and regularization. Applying Lemma 2.3, we get the following corollary.
@@
Under Assumption (2.3), denoting $\widehat{a}_{1}\geq\sup_{X^{\prime}\in\overline{\mathcal{X}},~{}KX^{\prime}=\mathbf{0}}\|(I-\Pi)\nabla F(X^{\prime})\|_{*}$ where $\nabla F:\overline{\mathcal{X}}\to\mathbb{R}^{md}$ is an arbitrary subgradient oracle, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v})$. If Assumption 2.1 also holds and $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$,
@@
Assume that all norms are $l_{2}$ norms, then with orthogonality, $KK^{*}$ is “diagonal” and so $\sigma_{\min}^{+}(K)=\min_{s\in[S]}\sigma_{\min}^{+}(K_{s})$. In addition, since $\|(I-\Pi)\nabla F(X)\|\geq\|\Pi_{s}\nabla F(X)\|$ for any $X\in\overline{\mathcal{X}}$ and $s\in[S]$, one can always take $a_{s}\leq\widehat{a}_{1}$ for all $s\in[S]$. Thus $\frac{\xi+\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}\geq\frac{\xi+a_{s}}{\sigma_{\min}^{+}(K_{s})}$. Since for the function $h(x)=\lambda\|x\|$ defined on $\mathbb{R}^{n_{0}}$ for some $\lambda>0$, the conjugate $h^{*}$ is the characteristic function of $\{x\in\mathbb{R}^{n_{0}}|\|x\|\leq\lambda\}$, i.e. $\mathsf{dom}(h^{*})=\{x\in\mathbb{R}^{n_{0}}|\|x\|\leq\lambda\}$. As will be seen in Theorem 3.1, the convergence is faster with smaller domains, suggesting that one should use the more refined decomposition when orthogonality holds.
@@
Comparisons with [lan_communication-efficient_2020] when $S=1$. Assume that $\widehat{R}^{prj}_{1}$ in (13) is used for some constant $\xi>0$ and $\widehat{a}_{1}=2\sqrt{m}M_{f}$, where $M_{f}$ (defined below) is an upper bound on the norm of the subgradient oracle $f^{\prime}_{v}\in\partial f_{v}$ (i.e. only one subgradient in the subdifferential for each $x\in\mathcal{X}$, $v\in V$). Then, the diameter of $\mathsf{dom}(R^{*})$ is $O(\frac{\sqrt{m}M_{f}}{\sigma^{+}_{\min}(K)})$. In [lan_communication-efficient_2020], it is shown that for ($\mathcal{P}_{s}$) with $R$ being the characteristic function of $\{\mathbf{0}\}$, there exists an optimal dual solution $\|Y^{*}\|\leq\frac{\sqrt{m}\widehat{M}_{f}}{\sigma^{+}_{\min}(K)}$, where $\widehat{M}_{f}$ is an upper bound on the norms of all subgradients $g\in\partial f_{v}$:
$\widehat{M}_{f}:=\sup_{x\in\mathcal{X},~{}v\in V,~{}g\in\partial f_{v}(x)}\|g\|_{*}\geq M_{f}:=\sup_{x\in\mathcal{X},~{}v\in V}\|f^{\prime}_{v}(x)\|_{*}.$
@@
Thus, even without function similarity, our $\widehat{R}^{prj}_{1}$ provides better control over the dual variables, leading to faster convergence. Moreover, when $\mathcal{X}\neq\mathbb{R}^{d}$, due to the normal cones at the boundary of $\mathcal{X}$, $\widehat{M}_{f}=\infty$.
3 Multi-timescale gradient sliding
To solve ($\mathcal{P}_{s}$) with the costs of information aggregation in mind, we resort to the decentralized communication sliding (DCS)[lan_communication-efficient_2020], a communication efficient variant of the classical primal-dual hybrid gradient algorithm. As a recap, at iteration $k$, DCS performs the following updates: ($R$ is the characteristic function of $\{\mathbf{0}\}$, and $w^{y}(Y)=\frac{1}{2}\langle Y,Y\rangle$)
$\widetilde{X}^{k}$ $=X^{k}+\alpha_{k}(X^{k}-X^{k-1})$ (14)
$Y^{k}$ $=\operatorname*{argmin}_{Y\in\mathbb{R}^{n}}R^{*}(Y)+\langle-K\widetilde{X}^{k},Y\rangle+\tau_{k}D_{w^{y}}(Y,Y^{k-1})$ (15)
$X^{k}$ $\approx\operatorname*{argmin}_{X\in\overline{\mathcal{X}}}F(X)+\langle K^{*}Y^{k},X\rangle+\eta_{k}D_{w^{X}}(X,X^{k-1})$ (16)
where the proximal update (16) is solved inexactly through the Communication-Sliding (CS) procedure (i.e. $T_{k}$ steps of mirror descent) locally by each primal agent. Since only computing $K\widetilde{X}^{k}$ and $K^{*}Y^{k}$ requires the communication between primal and dual agents, to reach $\epsilon$ suboptimality in terms of the gap, $O(1/\epsilon)$ rounds of communication/matrix-vector multiplication are needed. This is desirable, especially when the bottleneck of the entire algorithm is communication and/or matrix-vector multiplication.
We take a step further: with separable duals, we consider potential heterogeneity in the costs of information aggregation (e.g. sending messages and/or computing matrix-vector products). Since the dual (15) can be separated into $S$ blocks, one can apply block-coordinate descent type of algorithms such as S-PDHG [Chambolle2018Stochastic], which updates only a random subset of the dual blocks at each iteration. The flexibility in choosing the sampling distribution allows one to control the frequency of updates of different blocks. However, due to the randomness, the $O(1/k)$ rate of convergence is usually shown only for the expected gap.
To maintain the deterministic convergence guarantee as well as the flexibility in choosing the number of updates applied to each dual block, we propose the Multi-Timescale Gradient Sliding (MT-GS) for ($\mathcal{P}_{s}$), where different dual agents live in different timescales, and update at different rates: $\mathsf{Agent}(y_{s})$ only updates $y_{s}$ at iteration $0,r_{s},2r_{s},\ldots$. Our convergence results indicate that to reach $\epsilon$ duality gap, the number of communication rounds needed is $N=O(\frac{\overline{r}}{\epsilon})$, where $\overline{r}$ is the weighted average of $\{r_{s}\}_{s\in[S]}$, with weights depending on “dual domain sizes”, and thus on the function similarities if the penalties $R_{s}$’s are chosen as suggested by Section 2.3. In addition, when the costs of updating different duals and/or the function variation along different $K_{s}$’s are different (significantly), one can take advantage of the flexibility in choosing $\{r_{s}\}_{s\in[S]}$ to design more efficient algorithms.
3.1 Updating rules of multi-timescale gradient sliding
We assume that there is a global time $k=0,1,2,\ldots,N$.
Initialization. We assume that each $\mathsf{Agent}(x_{v})$ is given some $x_{v}^{init}\in\mathcal{X}$, and initialize $x_{v}^{k^{\prime}}=\widehat{x}_{v}^{k^{\prime}}=x_{v}^{init}$ for all $k^{\prime}<0$. Similarly, each $\mathsf{Agent}(y_{s})$ is given some $y_{s}^{init}\in\mathsf{dom}(R^{*}_{s})$, and initialize $y_{s}^{k^{\prime}}=y_{s}^{init}$ for all $k^{\prime}<0$.
Dual updates. We associate each dual $y_{s}$ with a rate $r_{s}\in\mathbb{N}$ and a local time $i_{s}=0,1,\ldots,N_{s}-1$, such that $N+1=r_{s}N_{s}$. For $k=0,1,\ldots,N$, $\mathsf{Agent}(y_{s})$ remains dormant (no computation or communication) unless $k=r_{s}i_{s}$ for some $i_{s}\in\{0,1,\ldots,N_{s}-1\}$, where $\mathsf{Agent}(y_{s})$ computes the updated $y^{i_{s}}_{s}$ using the following rules.
$\widetilde{x}_{s,v}^{i_{s}}$ $=\alpha_{s,i_{s}}\left(\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}}(\widehat{x}_{v}^{k^{\prime}}-x_{v}^{k^{\prime}-r_{s}})\right)+\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}+r_{s}}{x}_{v}^{k^{\prime}},\quad v\in V,$ (17)
$y_{s}^{i_{s}}$ $=\operatorname*{argmin}_{y_{s}\in\mathbb{R}^{n_{s}}}\langle-\frac{1}{\sum_{k^{\prime}=r_{s}i_{s}}^{r_{s}i_{s}+r_{s}-1}\theta_{k^{\prime}}}\sum_{v\in V}K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}\rangle+R^{*}_{s}(y_{s})+\tau_{s,i_{s}}D_{w^{y}_{s}}(y_{s},y^{i_{s}-1}_{s}).$ (18)
Further, we denote $\overline{y}_{s}^{k}=y_{s}^{\lfloor k/r_{s}\rfloor}$, i.e. the corresponding dual $y_{s}$ at the global time $k$, and $\overline{Y}^{k}=(\overline{y}_{s}^{k})_{s\in[S]}$
Primal updates. All primal variables are updated at each global time, through a generalized CS procedure, which we provide details in Section 3.2. For $k=0,1,\ldots,N$,
$(x_{v}^{k},\widehat{x}_{v}^{k})=CS(f_{v},{\mathcal{X}},D_{w^{x}},T_{k},(\eta_{k,s})_{s\in[S]},(x^{k-r_{s}}_{v})_{s\in[S]},\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},x_{v}^{k-1}),\quad\forall~{}v\in V.$ (19)
With correct choices of the parameters, (19) provides approximate solutions to the following problem
$\min_{x\in\mathcal{X}}f_{v}(x)+\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},x\rangle+\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(x,x_{v}^{k-r_{s}}).$ (20)
Algorithm 1 (Accelerated) Multi-timescale gradient sliding
$\{\alpha_{s,i_{s}}\},\{\theta_{k}\},\{T_{k}\},\{\eta_{k,s}\},\{\tau_{s,i_{s}}\},\{r_{s}\},X^{init},Y^{init}$
$(X^{k^{\prime}},\widehat{X}^{k^{\prime}},Y^{k^{\prime}})\leftarrow(X^{init},X^{init},Y^{init})$ for all $k^{\prime}<0$
for $k=0,1,\ldots,N$ $\triangleright$ implicitly $i_{s}=\lfloor k/r_{s}\rfloor$ for all $s\in[S]$
     for $s\in[S]$ such that $k=0~{}(\text{ mod }r_{s})$ do
@@ 3.1 Updating rules of multi-timescale gradient sliding
              $\mathsf{Agent}(x_{v})$ computes $\widetilde{x}_{s,v}^{i_{s}}$ using (17), and sends to $\mathsf{Agent}(y_{s})$
@@
         $\mathsf{Agent}(y_{s})$ computes $y_{s}^{i_{s}}$ using (18), and sends $y_{s}^{i_{s}}-y_{s}^{i_{s}-1}$ ($y_{s}^{0}$ if $i_{s}=0$) to $\mathsf{Agent}(x_{v})$ for all $v\in V$ such that $K_{s,v}\neq\mathbf{0}$
@@
         $\mathsf{Agent}(x_{v})$ computes $\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k}$, and updates $(x_{v}^{k},\widehat{x}_{v}^{k})$ using (19)
@@
Output $Z^{N}$ in (21).
Notice $\mathsf{Agent}(x_{v})$ needs to be able to compute all the $\widetilde{x}_{s,v}^{i_{s}}$, which could require extra space to store past $(x_{v},\widehat{x}_{v})$’s. One approach is to keep in memory all $x_{v}^{(k-2r_{\max}):(k-1)}$ and $\widehat{x}_{v}^{(k-r_{\max}):(k-1)}$, where $r_{\max}=\max_{s\in[S]}r_{s}$. This requires storing $3r_{\max}$ vectors in $\mathbb{R}^{d}$. Another approach is to keep in memory the accumulative $(x_{v},\widehat{x}_{v})$: for each $s\in[S]$, for $r_{s}(i_{s}-1)\leq k\leq r_{s}i_{s}$, keep in memory (previously computed) $\sum_{k^{\prime}=r_{s}i_{s}-2r_{s}}^{r_{s}i_{s}-r_{s}-1}\theta_{k^{\prime}+r_{s}}{x}_{v}^{k^{\prime}}$, while computing the sum $\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}}\widehat{x}_{v}^{k^{\prime}}$ and $\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\theta_{k^{\prime}+r_{s}}{x}_{v}^{k^{\prime}}$ as $k$ increases. This requires storing $3S$ vectors in $\mathbb{R}^{d}$. Also, due to the proximal centers, $x_{v}^{(k-r_{\max}):k}$ should be kept in memory, which requires storing $r_{\max}$ vectors.
In addition, notice that in Algorithm 1, $\mathsf{Agent}(y_{s})$ calculates $K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$ and sends a vector in $\mathbb{R}^{n_{s}}$, $\mathsf{Agent}(x_{v})$ calculates $K_{s,v}^{*}(\overline{y}_{s}^{k}-\overline{y}_{s}^{k-1})$ and sends a vector in $\mathbb{R}^{d}$. In fact, there are many task assignment strategies: for instance, $K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$ can also be computed by $\mathsf{Agent}(x_{v})$, and the message from $\mathsf{Agent}(x_{v})$ to $\mathsf{Agent}(y_{s})$ will be $K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$. This is preferable if $\mathsf{Agent}(x_{v})$ can compute matrix-vector products faster/at lower cost than $\mathsf{Agent}(y_{s})$. Due to this variability, in the cost analysis in Section 3.5, we take a “modular” perspective and assume that the cost of updating $y_{s}$ (including all matrix-vector multiplication and communication) is $c_{s}$.
3.2 Generalized communication sliding
In [lan_communication-efficient_2020], the CS procedure is used to approximately solve the primal proximal updates. However, due to differences in dual time scales, we need extra control on the variation of the primal sequence. Intuitively, when computing $X^{k}$, the $K_{s}^{*}y_{s}$ term is evaluated using $\overline{y}_{s}^{k}=y_{s}^{i_{s}}$, where $i_{s}=\lfloor k/r_{s}\rfloor$. However, $y_{s}^{i_{s}}$ is computed at the global iteration $r_{s}\lfloor k/r_{s}\rfloor\leq k$, using the (outdated) $X_{s}^{(r_{s}i_{s}-2r_{s}):(r_{s}i_{s}-1)}$ and $\widehat{X}_{s}^{(r_{s}i_{s}-r_{s}):(r_{s}i_{s}-1)}$. Thus, for the multi-timescale updates to converge, we need to control the variation of the primal sequence. To achieve that, we take the proximal term in the primal updates as a mixture of proximal terms centered at $\{X^{k-r_{s}}\}_{s\in[S]}$, which gives the formulation (20).
Motivated by this, we propose the following generalization of the CS procedure in [lan_communication-efficient_2020] (which is a special case where $|\mathcal{I}|=1$).
Algorithm 2 Generalized communication sliding procedure
$(u^{T},\widehat{u}^{T})=CS(\phi,U,D_{w},T,(\eta_{i})_{i\in\mathcal{I}},v,(x_{i})_{i\in\mathcal{I}},x^{init})$, an approximate solution to
$\min_{u\in U}\Phi(u):=\langle v,u\rangle+\phi(u)+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i})$
for $t=1,\ldots,T$ do
$u^{t}=\operatorname*{argmin}_{u\in U}\langle v+\phi^{\prime}(u^{t-1}),u\rangle+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i})+\eta\beta_{t}D_{w}(u,u^{t-1})$
end for
$\widehat{u}^{T}=(\sum_{t=1}^{T}\lambda_{t})^{-1}\sum_{t=1}^{T}\lambda_{t}u^{t}.$
As a corollary to Lemma B.2, we have the following performance guarantee.

### v2

@@ 4.3.2 Function similarity for general convex functions
Assume that for all $s\neq s^{\prime}\in[S]$, $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$. We say that the set of functions $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar if there exists a subgradient oracle $\nabla F:{\mathcal{X}}\to\mathbb{R}^{d}$, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v}))$, such that for each $s\in[S]$,
@@
$a_{1}^{2}\geq\sup_{x\in\overline{\mathcal{X}}}\sum_{v\in V}\|f^{\prime}_{v}(x)-\frac{1}{m}\sum_{v^{\prime}\in V}f_{v^{\prime}}^{\prime}(x)\|^{2}.$
@@
Comparisons with existing notions of function similarity. [22] proposes the bounded gradient dissimilarity for differentiable convex objectives, which coincides with our Definition 4.1 when $S=1$ and when the objectives are differentiable. For twice differentiable objectives, function similarity is also defined in terms of differences in Hessians, i.e. $\|\nabla^{2}f_{v}-\nabla^{2}f_{v^{\prime}}\|$ [48, 24, 4, 22]. For general convex functions which could be non-differentiable, [4] informally defines it ($\delta$-relatedness in their terminology) as the condition that “subgradients of local functions are at most $\delta$-different from each other”. Our Definition 4.1 formalizes this idea, and extend it to the case where $S>1$.
@@ 4.3.3 Requirements on $R_{s}$ without orthogonality
The above Lemma 4.3 imposes orthogonality assumptions on $\{K_{s}\}_{s\in[S]}$. In the more general case where such assumptions do not hold, one can always view ($\mathcal{P}_{b}$) as a problem with only $1$ block, with $K$ and $R$ as the corresponding operator and regularization. Applying Lemma 4.3, we get the following corollary.
@@
Denoting $\widehat{a}_{1}\geq\sup_{X^{\prime}\in{\mathcal{X}},~KX^{\prime}=\mathbf{0}}\|(I-\Pi)\nabla F(X^{\prime})\|_{*}$ where $\nabla F:{\mathcal{X}}\to\mathbb{R}^{d}$ is an arbitrary subgradient oracle, i.e. $(\nabla F(X))_{v}\in\partial f_{v}(x_{v})$. If $\sup_{Y^{\prime}\in\mathsf{dom}(R^{*})}G(X,Y;X^{*},Y^{\prime})\leq\epsilon$,
@@
Comparisons with [27] when $S=1$. Assume that $\widehat{R}^{prj}_{1}$ in (28) is used for some constant $\xi>0$ and $\widehat{a}_{1}=2\sqrt{m}M_{f}$, where $M_{f}$ (defined below) is an upper bound on the norm of the subgradient oracle $f^{\prime}_{v}\in\partial f_{v}$ (i.e. only one subgradient in the subdifferential for each $x\in\overline{\mathcal{X}}$, $v\in V$). Then, the diameter of $\mathsf{dom}(R^{*})$ is $O(\frac{\sqrt{m}M_{f}}{\sigma^{+}_{\min}(K)})$. In [27], it is shown that for ($\mathcal{P}_{b}$) with $R$ being the characteristic function of $\{\mathbf{0}\}$, there exists an optimal dual solution $\|Y^{*}\|\leq\frac{\sqrt{m}\widehat{M}_{f}}{\sigma^{+}_{\min}(K)}$, where $\widehat{M}_{f}$ is an upper bound on the norms of all subgradients $g\in\partial f_{v}$:
$\widehat{M}_{f}:=\sup_{x\in\overline{\mathcal{X}},~v\in V,~g\in\partial f_{v}(x)}\|g\|_{*}\geq M_{f}:=\sup_{x\in\overline{\mathcal{X}},~v\in V}\|f^{\prime}_{v}(x)\|_{*}.$
@@
Thus, even without function similarity, our $\widehat{R}^{prj}_{1}$ provides better control over the dual variables, leading to faster convergence.
4.4 Applying (accelerated) multi-timescale PDHG to distributed optimization
In this section, we present the distributed implementation of Algorithm 1 to the primal-dual formulation of the problem ($\mathcal{P}_{d}^{lift}$). We first point it out that if the distance generating function $w_{X}$ in the lifted space is separable, i.e. $w_{X}(X)=\sum_{v\in V}w_{x_{v}}(x_{v})$, then the Bregman divergence is also separable, i.e. $D(X,X^{\prime})=\sum_{v\in V}D(x_{v},x^{\prime}_{v})$. Now revising the updates (3c) and (3a) as well as our proposed multi-timescale updates (5a), (6), and the approximation using the generalized communication sliding (8), we see that they are all decomposable w.r.t. the primal agents, and can be implemented locally by each $\mathsf{Agent}(x_{v})$ without any communication. More precisely, defining
$\phi_{v}^{k}(x_{v}):=f_{v}(x_{v})+\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},x_{v}\rangle+\sum_{s=1}^{S}\eta_{k,s}D(x_{v},x_{v}^{k-r_{s}}),$ (29)
then we require that
$\phi_{v}^{k}(\widehat{x}_{v}^{k})\leq\phi_{v}^{k}(x_{v})-(\frac{\mu_{f}}{C}+\eta_{k})D(x_{v},x_{v}^{k})+\delta_{k,v}(x_{v}),\quad\forall x_{v}\in\overline{\mathcal{X}},$ (30)
which can be achieved through the generalized communication sliding procedure applied to each $x_{v}$ locally by $\mathsf{Agent}(x_{v})$. Thus, the only communication needed for the primal dual updates is to make sure the following two conditions are met:
  • the dual $\mathsf{Agent}(y_{s})$ knows $K_{s}\widetilde{X}_{s}^{i_{s}}$ at iteration $k=r_{s}i_{s}$, which can be realized if at the beginning of iteration $k=r_{s}i_{s}$, each primal agent $v$ computes $\widetilde{x}_{v}^{i_{s}}$ and send it to the dual agent $y_{s}$, then the dual agent computes $\sum_{v\in V}K_{s,v}\widetilde{x}_{v}^{i_{s}}$;
  • the primal $\mathsf{Agent}(x_{v})$ knows $\sum_{s=1}^{S}K^{*}_{s,v}\overline{y}_{s}^{k}$ at iteration $k$; this can be achieved if after each dual update, the dual agent $y_{s}$ sends $y_{s}^{i_{s}}-y_{s}^{i_{s}-1}$ to all primal agents, then the primal agents can update $\sum_{s=1}^{S}K^{*}_{s,v}\overline{y}_{s}^{k}$ using $K^{*}_{s,v}(y_{s}^{i_{s}}-y_{s}^{i_{s}-1})$.
We provide such implementation in Algorithm 2, and explicitly mark the steps which require communication in green. Here, we provide few remarks. Notice that in Algorithm 1, $\mathsf{Agent}(y_{s})$ calculates $\sum_{v\in V}K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$ and sends $y_{s}^{i_{s}}-y_{s}^{i_{s}-1}$, a vector in $\mathbb{R}^{n_{s}}$, $\mathsf{Agent}(x_{v})$ calculates $K_{s,v}^{*}(\overline{y}_{s}^{k}-\overline{y}_{s}^{k-1})$ and sends $\widetilde{x}_{s,v}^{i_{s}}$, a vector in $\mathbb{R}^{\overline{d}}$. In fact, there are many task assignment strategies: for instance, $K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$ can also be computed by $\mathsf{Agent}(x_{v})$, and the message from $\mathsf{Agent}(x_{v})$ to $\mathsf{Agent}(y_{s})$ will be $K_{s,v}\widetilde{x}_{s,v}^{i_{s}}$. This is preferable if $\mathsf{Agent}(x_{v})$ can compute matrix-vector products faster/at lower cost than $\mathsf{Agent}(y_{s})$. Due to this variability, in the cost analysis below, we take a “modular” perspective and assume that the cost of updating $y_{s}$ (including all matrix-vector multiplication and communication) is $c_{s}$.
Algorithm 2 (Accelerated) Multi-timescale PDHG for distributed optimization
$\{\alpha_{s,i_{s}}\},\{\theta_{k}\},\{\eta_{k,s}\},\{\tau_{s,i_{s}}\},\{r_{s}\},X^{init},Y^{init}$
Primal dual pair $Z^{N}$
Initialize $(X^{k^{\prime}},\widehat{X}^{k^{\prime}},Y^{k^{\prime}})\leftarrow(X^{init},X^{init},Y^{init})$ for all $k^{\prime}<0$
[…]
@@ 4.4 Applying (accelerated) multi-timescale PDHG to distributed optimization
     $\mathsf{Agent}(x_{v})$ computes $\widetilde{X}_{s}^{i_{s}}$ using (5a) and sends it to $\mathsf{Agent}(y_{s})$
@@
   Dual update: $\mathsf{Agent}(y_{s})$ computes $\sum_{v\in V}K_{s,v}\widetilde{x}_{v}^{i_{s}}$, then updates $y_{s}^{i_{s}}$ using (5b)
   $\mathsf{Agent}(y_{s})$ sends $y_{s}^{i_{s}}-y_{s}^{i_{s}-1}$ ($y_{s}^{0}$ if $i_{s}=0$) to $\mathsf{Agent}(x_{v})$ for all $v\in V$ such that $K_{s,v}\neq\mathbf{0}$
@@
   Primal update: $\mathsf{Agent}(x_{v})$ computes $K^{*}\overline{Y}^{k}$ where $\overline{y}_{s}^{k}=y_{s}^{\lfloor k/r_{s}\rfloor}$ for $s\in[S]$, updates $(x_{v}^{k},\widehat{x}_{v}^{k})$ satisfying (30)
@@
All $\mathsf{Agent}(x_{v})$ and $\mathsf{Agent}(y_{s})$ computes their components of $Z^{N}$ using (9).
4.5 Convergence for general convex objectives
With additional assumptions specific to distributed optimization, and with proper choices of $R_{s}$’s, the duality gap for ($\mathcal{P}_{d}^{lift}$) can be related to the suboptimality in terms of objective values $F$ and/or violation of the consensus constraint for the original problem ($\mathcal{P}_{d}$). Next, we establish such connection.